# Low-Altitude Wireless Networks: A Comprehensive Survey

### v2

@@ 4.3 Mobile Edge Computing
Mobile edge computing (MEC) seamlessly integrates computational resources into the network edge [li2025edge], enabling data processing and inference to occur locally at edge nodes such as UAVs, base stations, and edge servers. By minimizing the need to transmit large volumes of raw data to distant cloud data centers, this approach effectively mitigates latency and alleviates bandwidth congestion, thereby enhancing system efficiency, scalability, and resilience.
@@
This progression is especially pivotal for enabling and sustaining LAE activities, where UAV operations must continuously adapt to dynamic and uncertain environments. LAE-relevant tasks, such as real-time logistics routing, aerial infrastructure inspection, target tracking, and environmental perception, require low-latency and highly reliable computation. Traditional cloud-centric computing models are often constrained by high latency due to long-distance backhaul links, which can hinder time-sensitive applications. In contrast, MEC enables lightweight computational models to be deployed directly on UAVs or adjacent edge nodes, facilitating immediate decision-making and significantly enhancing responsiveness, autonomy, and mission reliability [huynh2022uav]. This results in more efficient and reliable execution across a diverse array of low-altitude economic scenarios.
Furthermore, the distributed nature of LAWNs, with multiple UAVs operating cooperatively, lends itself well to distributed computing paradigms such as federated learning. These paradigms offer scalable, privacy-preserving, and resource-efficient solutions for LAE operations. By training models collaboratively across UAV nodes without sharing raw data, federated learning not only ensures privacy but also reduces communication overhead [zhang2025latency]. For example, [deng2022uav] introduced a UAV-enabled MEC framework for air-ground integrated networks, where UAVs acted as aerial base stations to provide adaptive coverage and edge computing support for Internet-of-Things (IoT) devices under stringent energy and accuracy constraints. Similarly, [wang2023uav] proposed a two-tier hierarchical federated learning framework, where UAVs served as mobile relays, enhancing connectivity to a central server while simultaneously conducting local computations, thus improving scalability in large-scale, infrastructure-limited environments.
Beyond distributed learning, recent advancements in UAV-enabled MEC have explored a variety of application domains and performance metrics directly applicable to LAE services. For instance, [zhang2025age] introduced a UAV-assisted IoT data collection framework under segmented fading channels, which achieved reduced AoI, lower energy consumption, and faster mission completion times through collaborative UAV operations. In [shaikh2025flying], a flying-edge computing framework for smart agriculture was proposed, where UAVs equipped with AI-powered sensors perform tasks such as disease detection and irrigation control on-board, enabling real-time decision-making in areas with limited connectivity. Additionally, [zhang2024minimizing] presented a UAV-assisted MEC framework that optimizes UAV deployment and computation offloading to minimize response delays, ensuring rapid adaptability to dynamic operational environments.
@@ 4.4 Wireless Power Transfer
Wireless power transfer (WPT) delivers energy to electronic devices through radio-frequency electromagnetic waves or inductive coupling, thereby enabling continuous power replenishment without physical connections [zhang2018wireless]. In LAWNs, WPT provides an effective solution to the endurance limitations of battery-powered UAVs, extending mission duration, reducing task interruptions, and minimizing maintenance costs [xu2018uav]. Its integration with UAV swarms further enhances operational sustainability, making long-duration and high-frequency missions feasible. The evolution of WPT in UAV-enabled networks can be conceptually divided into three progressive stages, as illustrated in Fig. 5.
@@
Extending the concept of cooperative energy sharing, recent research has explored the implementation of LAWNs-enabled large-scale WPT in practical scenarios, which can serve as a key technological enabler for various LAE applications. In [hu2020wireless], a cooperative wireless-powered MEC framework was proposed, where a grid-powered access point provided both computational services and laser-based charging to a UAV. The UAV, in turn, functioned as an information relay, energy relay, and mobile edge server. Similarly, [yin2019uav] introduced a UAV-assisted cooperative system that employed simultaneous wireless information and power transfer (SWIPT), enabling UAVs to act as mobile relays powered directly by RF energy. Beyond cooperative relaying, heterogeneous architectures were explored. For instance, [liu2024joint] proposed a UAV-enabled MEC framework where UAVs served as both wireless chargers and mobile computing enhancers for energy-constrained IoT devices. Likewise, [xie2018throughput] considered a harvest-then-transmit protocol in UAV-enabled wireless powered communication networks, while [wang2025practical] investigated trajectory optimization in obstacle-rich environments to maximize energy delivery efficiency. At the ground-user level, [li2023energy] optimized 3D UAV trajectories for fair and efficient power allocation among distributed nodes, and [shi2023two] addressed charging of energy-constrained sensor nodes with unknown locations via UAV-based WPT strategies. In summary, WPT in LAWNs evolves from static ground-based charging to a dynamic, network-wide energy sharing paradigm. By leveraging UAV mobility, swarm cooperation, and joint optimization with sensing and communication, LAWNs enable persistent, autonomous, and large-scale aerial operations. When combined with ISAC and edge intelligence, WPT does not merely extend flight time. It provides the energy backbone that sustains intelligent perception, resilient communication, and adaptive networked control for low-altitude economic scenarios such as long-endurance surveillance, infrastructure inspection, cargo transport, and emergency energy replenishment.
@@ 5.1 Privacy preservation
Privacy preservation of LAWNs refers to protecting sensitive information and preventing unauthorized access or disclosure of data related to the operation, environment, and participants within LAWNs [cordill2025comprehensive, kaleem2025quantum]. By addressing the privacy concerns of the above issues, LAWNs can foster user trust, enable wider adoption, and comply with legal and ethical requirements, ultimately unlocking the full potential of LAWN technology while minimizing the risks associated with data breaches and misuse. Currently, researchers have adopted differential privacy, federated learning, homomorphic encryption, secure multi-party computation, destination obfuscation, game theory and blockchain-based techniques to realize the privacy preservation of LAWNs on sensing, communication, computing, and control aspects.
@@
For example, for privacy preservation on the sensing aspect, to ensure privacy in UAV-based delivery and sensing, [liu2023decentralized] introduced a game-theoretic routing framework that aims to protect UAV trajectory privacy. The approach leverages a task-time graph and a non-cooperative potential game to optimize sensing, routing, and task selection. This framework demonstrates effective privacy preservation while ensuring timely delivery and achieving efficiency. In terms of communication privacy, [chen2020traceable] presented a traceable and privacy-preserving authentication scheme for UAV applications, employing digital signatures to ensure integrity, confidentiality, and manage sensitive area access. Addressing security in UAV surveillance, [deebak2020smart] introduced a smart Internet of drone (S-IoD) framework with IPA and a lightweight privacy-preserving scheme (L-PPS). This L-PPS expedites authentication between devices through a secret token and dynamic user authentication, mitigating computational overhead. For privacy in UAV-assisted IoT computation offloading, [wei2021uav] identified that current deep reinforcement learning-based methods leak UAV offloading preferences. The authors introduced a privacy-preserving method that integrates differential privacy into deep reinforcement learning to protect these preferences, including formal analysis of security and utility. Similarly, as shown in Fig. 6, [wang2023seal] addressed the critical requirements of incentives, privacy, and fairness when offloading UAV computation tasks to ground vehicles. Their proposed framework overcomes issues of manipulation and privacy leakage by using a strategy-proof reverse combinatorial auction. In addition, fairness is managed through smart contracts and hash-chain micropayments, complemented by a privacy-preserving off-chain auction protocol utilizing a trusted processor. Focusing on control-related privacy, [pan2019unmanned] addressed the privacy-preserving navigation (PPN) problem for visual-based UAVs, aiming to minimize privacy violations within specific sub-areas while covering a target region through a heuristic path planning solution. The work of [liu2024privacy] tackled privacy concerns in cooperative UAV operations by proposing a decentralized framework that preserves trajectory and destination privacy through probabilistic destination obfuscation and a noncooperative Bayesian game for joint routing and charging.
@@
Figure 6: The framework of the proposed SEAL defense scheme in UAV offloading scenario [wang2023seal].
@@ 5.2 Physical-Layer security
Physical layer security of LAWNs refers to protecting the confidentiality, integrity, and availability of information transmitted and processed within LAWNs by exploiting the inherent characteristics of the physical layer itself [wang2021physical, 10168298]. This protection is achieved without relying solely on traditional cryptographic methods. By addressing the physical layer security concerns of the above issues, LAWNs can enhance resilience against sophisticated attacks, enable robust operation in contested environments, and provide lightweight security solutions suitable for resource-constrained UAV platforms, ultimately contributing to safer and more reliable UAV deployments. In general, physical layer security is commonly classified into two fundamental paradigms, namely the wiretap channel and covert communication. Currently, researchers have adopted beamforming techniques, artificial noise injection, cooperative jamming, power allocation strategies, and intelligent reflecting surfaces (IRS) to realize the physical layer security of LAWNs on sensing, communication, computing, and control aspects. For example, for physical layer security in integrated sensing and communication, [xiu2024improving] explored a system where a UAV transmits ISAC waveforms to securely communicate with IoT devices while simultaneously sensing the environment. This approach optimizes UAV trajectory and beamforming to maximize the average communication secrecy rate while ensuring accurate sensing capabilities. In the context of covert ISAC in LAWN systems, [xx] studied covert ISAC for multi-antenna UAV systems using a unified waveform for communication and sensing, revealing a trade-off between covert communication and sensing accuracy by exploiting prior information of the adversary. Regarding physical layer security in UAV communication, [li2021physical] addressed wire-tap channel communications in 5G networks with eavesdroppers by proposing a virtual antenna array with collaborative beamforming. To enhance security, UAV positions and excitation current weights are optimized, aiming to improve total secrecy rates and reduce sidelobe levels through an improved multi-objective dragonfly algorithm. Furthermore, to satisfy the heterogeneous security requirements in low-altitude UAV systems, where some users demand secrecy while others require covertness, the authors in [10681761] investigated multi-user cooperative secret and covert communications for UAV networks. For the physical layer security aspect within computing, [zhou2019secure] investigated a secure UAV MEC system where ground users offload tasks to a legitimate UAV amidst eavesdropping UAVs. It enhances security by employing jamming signals from the legitimate UAV and users to maximize minimum secrecy capacity, jointly optimizing aspects like offloading ratio and computing capacity to balance security and latency. [nasir2024secure] focused on secure and energy-efficient UAV-MEC with RIS assistance, aiming to maximize minimum energy efficiency against an eavesdropper. The research ensured security through artificial noise transmission and optimizes computation tasks/CPU frequency allocation alongside other parameters, demonstrating strong performance even with an eavesdropper present. Focusing on physical layer security in control, [xu2025collaborative] proposed a secret and covert transmission strategy for ground-to-UAV links using non-orthogonal multiple access (NOMA). This method allows simultaneous transmissions with different security needs, enabling an adversary to attempt both data eavesdropping and transmission detection. Security is rigorously ensured with optimization targeting maximized rates through beamforming and trajectory control. [chen2021uav] improved covert communication range by using a UAV relay and finite blocklength transmissions, aiming to maximize secure bits against a warden. It achieves this by optimizing the warden’s threat potential and jointly controlling transmission parameters under error constraints, leveraging signal uncertainty at the warden for physical layer covertness.
@@ 5.3 Network-Layer security
For example, addressing network layer security challenges for UAV-based sensing, [sedjelmaci2017hierarchical] designed an intrusion detection and response system to identify and mitigate attacks that compromise the integrity of sensed data, including false information dissemination and GPS spoofing. To ensure network layer security in UAV communication, [fotohi2022self] introduced self-adaptive intrusion detection (SID)-UAV, which employs a self-matching system to detect safe communication routes, coupled with a distributed architecture for investigation and response to malicious UAVs. The work of [mao2024energy] enhanced network layer security for autonomous aerial vehicle (AAV)-aided short-packet transmissions against detection and eavesdropping by optimizing for maximum covert secrecy energy efficiency through a dual-loop iterative algorithm for user association, power control, and trajectory planning. Addressing network layer security in UAV edge computing, [sedjelmaci2019efficient] presented a cyber-defense mechanism based on a non-cooperative game to protect against network and offloading threats. The solution aims to maximize protection while minimizing the computational and energy costs associated with the defense. In response to the data leakage risk in UAV-assisted traffic surveillance at the network layer, [garg2018uav] utilized a probabilistic data structure approach for cyber-threat detection and efficient data management, minimizing computational overhead in the process. Regarding network layer security for control, [ale2024advanced] proposed using sliding mode control to enhance UAV security. The system dynamically reconfigures to adapt to network-level threats like GPS spoofing and jamming, ensuring robust control even under attack through anomaly detection and fault tolerance. [xie2025blockchain] tackled network layer access control challenges in distributed UAV swarm control by using a blockchain-based scheme. It ensures secure authentication and data integrity against identity spoofing and data injection, employing dynamic key management for task and identity groups.
@@ 6.1 Large AI Model-based LAWNs
Large AI models, characterized by their extensive parameter dimensions and training on massive datasets, have made notable advancements across various domains, including natural language processing, computer vision, and autonomous systems. Prominent models, e.g., generative pre-trained transformers (GPT), bidirectional encoder representations from transformers (BERT), and similar architectures, possess the ability to learn intricate patterns and generalize across a wide range of tasks. This makes them particularly suitable for dynamic and complex environments. Hence, the integration of large AI models into LAWNs holds substantial potential for enhancing the cognitive capabilities of UAVs, facilitating real-time decision-making in unpredictable environments. A common approach to this integration involves utilizing large language models (LLMs) in physical layer designs, where pre-trained models are deployed at both the UAVs and controller terminal to enable effective online inference. For example, the work of [Xiao2024] illustrates how large AI models can be employed in UAV path planning, where they enable the UAVs to dynamically adjust their flight paths based on real-time environmental data, such as wind and obstacle conditions. Similarly, the work of [Javaid2024] explores how large language models (LLMs) can enhance UAV communication by enabling more intuitive interactions between UAVs and ground control stations. This capability allows for more adaptive and flexible communication, improving operational coordination, especially in complex scenarios like disaster relief or surveillance operations.
@@
Model distribution involves partitioning a large AI model across multiple devices or computational nodes, such as UAVs or edge servers. Instead of running the entire model on a single device, each UAV or edge server processes a portion of the model. This distributed approach allows for parallel processing, which significantly reduces the computational burden on each individual device. The work of [Piggott2023] demonstrated how model distribution can be effectively implemented in UAV networks, where large AI models are divided across several UAVs, enabling real-time inference without overwhelming the resources of any single UAV. This technique enhances the scalability of large models, making it possible to deploy them in a distributed manner without compromising system performance. Data distribution, on the other hand, focuses on distributing the data processing tasks across multiple devices to minimize the need for large-scale data transmission. In the context of LAWNs, this approach ensures that data collected by UAVs is processed locally, reducing the communication overhead and enabling real-time decision-making. The work of [Javaid2024] highlighted the use of federated learning in UAV networks, a data distribution method that allows multiple UAVs to collaboratively train a shared model by exchanging model updates rather than raw data. This approach not only preserves data privacy but also reduces the bandwidth required for communication between UAVs and ground control systems, making it particularly beneficial in applications where security and privacy are paramount.
@@
However, it is important to note that both model distribution and data distribution entail substantial resource-intensive fine-tuning processes [10415235]. Therefore, high-dimensional computing (HDC) has attracted increasing attention as a computationally efficient alternative [chang2023recent]. HDC is a brain-inspired computing paradigm in which information is encoded into extremely high-dimensional vectors, often containing tens of thousands of components. Such representations naturally support large-scale parallelism and enable rapid processing, while exhibiting strong robustness against noise, distortions, and missing data. A fundamental property of these high-dimensional vectors is that independently generated vectors are almost orthogonal to one another. This behavior results from the randomization procedures during vector construction, which ensure that distinct vectors share very limited overlap [basaklar2021hypervector]. Consequently, meaningful representations tend to have large Hamming distances and low cosine similarity, providing substantial capacity for encoding diverse information with minimal interference. These characteristics enable HDC to rely on simple arithmetic operations, including addition for combining multiple vectors and multiplication for binding variables with their associated values [kanerva2009hyperdimensional]. Similarity measures, such as Hamming distance and cosine similarity, are used to determine how closely two vectors align, supporting accurate retrieval and comparison. Since these operations do not require heavy matrix computations, they are highly efficient and particularly suitable for UAVs operating in LAWNs, where memory, computation capability, and energy supply are inherently limited.
@@
HDC has also shown strong potential for improving wireless communication reliability in scenarios relevant to LAWNs. Its high-dimensional representations can tolerate interference and signal distortions, allowing communication to remain stable even when multiple devices transmit simultaneously. A representative example is the direct transmission of high-dimensional vectors from a group of sensing devices. Instead of coordinating access or assigning dedicated channels, each device simply sends the vector that corresponds to its measurement. The receiver then observes the superimposed vector and determines the underlying sensing events by comparing the received vector with a set of reference patterns. By doing so, HDC enables efficient UAV reporting without traditional access control. For example, the authors in [jakimovski2012collective] considered a scenario where the receiver detects temperature conditions and determines how many devices experienced a particular temperature by analyzing the combined vector. Furthermore, the study in [kleyko2012dependable] generated composite vectors through binding and superposition to encode local sensing information at each device. The receiver successfully interpreted the embedded information using predefined base vectors, demonstrating that HDC can maintain reliable information extraction even in dense and interference-prone environments. These results indicate that HDC offers a practical and lightweight computational method for LAWNs, enabling robust communication and efficient information exchange while accommodating the limitations of UAV platforms.

### v3

@@ 4.3 Mobile Edge Computing
Mobile edge computing (MEC) seamlessly integrates computational resources into the network edge [94], enabling data processing and inference to occur locally at edge nodes such as UAVs, base stations, and edge servers. By minimizing the need to transmit large volumes of raw data to distant cloud data centers, this approach effectively mitigates latency and alleviates bandwidth congestion, thereby enhancing system efficiency, scalability, and resilience.
@@
This progression is especially pivotal for enabling and sustaining LAE activities, where UAV operations must continuously adapt to dynamic and uncertain environments. LAE-relevant tasks, such as real-time logistics routing, aerial infrastructure inspection, target tracking, and environmental perception, require low-latency and highly reliable computation. Traditional cloud-centric computing models are often constrained by high latency due to long-distance backhaul links, which can hinder time-sensitive applications. In contrast, MEC enables lightweight computational models to be deployed directly on UAVs or adjacent edge nodes, facilitating immediate decision-making and significantly enhancing responsiveness, autonomy, and mission reliability [66]. This results in more efficient and reliable execution across a diverse array of low-altitude economic scenarios.
Furthermore, the distributed nature of LAWNs, with multiple UAVs operating cooperatively, lends itself well to distributed computing paradigms such as federated learning. These paradigms offer scalable, privacy-preserving, and resource-efficient solutions for LAE operations. By training models collaboratively across UAV nodes without sharing raw data, federated learning not only ensures privacy but also reduces communication overhead [200]. For example, [44] introduced a UAV-enabled MEC framework for air-ground integrated networks, where UAVs acted as aerial base stations to provide adaptive coverage and edge computing support for Internet-of-Things (IoT) devices under stringent energy and accuracy constraints. Similarly, [156] proposed a two-tier hierarchical federated learning framework, where UAVs served as mobile relays, enhancing connectivity to a central server while simultaneously conducting local computations, thus improving scalability in large-scale, infrastructure-limited environments.
Beyond distributed learning, recent advancements in UAV-enabled MEC have explored a variety of application domains and performance metrics directly applicable to LAE services. For instance, [201] introduced a UAV-assisted IoT data collection framework under segmented fading channels, which achieved reduced AoI, lower energy consumption, and faster mission completion times through collaborative UAV operations. In [142], a flying-edge computing framework for smart agriculture was proposed, where UAVs equipped with AI-powered sensors perform tasks such as disease detection and irrigation control on-board, enabling real-time decision-making in areas with limited connectivity. Additionally, [192] presented a UAV-assisted MEC framework that optimizes UAV deployment and computation offloading to minimize response delays, ensuring rapid adaptability to dynamic operational environments.
@@ 4.4 Wireless Power Transfer
Wireless power transfer (WPT) delivers energy to electronic devices through radio-frequency electromagnetic waves or inductive coupling, thereby enabling continuous power replenishment without physical connections [202]. In LAWNs, WPT provides an effective solution to the endurance limitations of battery-powered UAVs, extending mission duration, reducing task interruptions, and minimizing maintenance costs [177]. Its integration with UAV swarms further enhances operational sustainability, making long-duration and high-frequency missions feasible. The evolution of WPT in UAV-enabled networks can be conceptually divided into three progressive stages, as illustrated in Fig. 5.
@@
Extending the concept of cooperative energy sharing, recent research has explored the implementation of LAWNs-enabled large-scale WPT in practical scenarios, which can serve as a key technological enabler for various LAE applications. In [64], a cooperative wireless-powered MEC framework was proposed, where a grid-powered access point provided both computational services and laser-based charging to a UAV. The UAV, in turn, functioned as an information relay, energy relay, and mobile edge server. Similarly, [185] introduced a UAV-assisted cooperative system that employed simultaneous wireless information and power transfer (SWIPT), enabling UAVs to act as mobile relays powered directly by RF energy. Beyond cooperative relaying, heterogeneous architectures were explored. For instance, [105] proposed a UAV-enabled MEC framework where UAVs served as both wireless chargers and mobile computing enhancers for energy-constrained IoT devices. Likewise, [173] considered a harvest-then-transmit protocol in UAV-enabled wireless powered communication networks, while [159] investigated trajectory optimization in obstacle-rich environments to maximize energy delivery efficiency. At the ground-user level, [92] optimized 3D UAV trajectories for fair and efficient power allocation among distributed nodes, and [144] addressed charging of energy-constrained sensor nodes with unknown locations via UAV-based WPT strategies. In summary, WPT in LAWNs evolves from static ground-based charging to a dynamic, network-wide energy sharing paradigm. By leveraging UAV mobility, swarm cooperation, and joint optimization with sensing and communication, LAWNs enable persistent, autonomous, and large-scale aerial operations. When combined with ISAC and edge intelligence, WPT does not merely extend flight time. It provides the energy backbone that sustains intelligent perception, resilient communication, and adaptive networked control for low-altitude economic scenarios such as long-endurance surveillance, infrastructure inspection, cargo transport, and emergency energy replenishment.
@@ 5.1 Privacy preservation
Privacy preservation of LAWNs refers to protecting sensitive information and preventing unauthorized access or disclosure of data related to the operation, environment, and participants within LAWNs [36, 79]. By addressing the privacy concerns of the above issues, LAWNs can foster user trust, enable wider adoption, and comply with legal and ethical requirements, ultimately unlocking the full potential of LAWN technology while minimizing the risks associated with data breaches and misuse. Currently, researchers have adopted differential privacy, federated learning, homomorphic encryption, secure multi-party computation, destination obfuscation, game theory and blockchain-based techniques to realize the privacy preservation of LAWNs on sensing, communication, computing, and control aspects.
@@
For example, for privacy preservation on the sensing aspect, to ensure privacy in UAV-based delivery and sensing, [101] introduced a game-theoretic routing framework that aims to protect UAV trajectory privacy. The approach leverages a task-time graph and a non-cooperative potential game to optimize sensing, routing, and task selection. This framework demonstrates effective privacy preservation while ensuring timely delivery and achieving efficiency. In terms of communication privacy, [30] presented a traceable and privacy-preserving authentication scheme for UAV applications, employing digital signatures to ensure integrity, confidentiality, and manage sensitive area access. Addressing security in UAV surveillance, [43] introduced a smart Internet of drone (S-IoD) framework with IPA and a lightweight privacy-preserving scheme (L-PPS). This L-PPS expedites authentication between devices through a secret token and dynamic user authentication, mitigating computational overhead. For privacy in UAV-assisted IoT computation offloading, [163] identified that current deep reinforcement learning-based methods leak UAV offloading preferences. The authors introduced a privacy-preserving method that integrates differential privacy into deep reinforcement learning to protect these preferences, including formal analysis of security and utility. Similarly, as shown in Fig. 6, [160] addressed the critical requirements of incentives, privacy, and fairness when offloading UAV computation tasks to ground vehicles. Their proposed framework overcomes issues of manipulation and privacy leakage by using a strategy-proof reverse combinatorial auction. In addition, fairness is managed through smart contracts and hash-chain micropayments, complemented by a privacy-preserving off-chain auction protocol utilizing a trusted processor. Focusing on control-related privacy, [125] addressed the privacy-preserving navigation (PPN) problem for visual-based UAVs, aiming to minimize privacy violations within specific sub-areas while covering a target region through a heuristic path planning solution. The work of [102] tackled privacy concerns in cooperative UAV operations by proposing a decentralized framework that preserves trajectory and destination privacy through probabilistic destination obfuscation and a noncooperative Bayesian game for joint routing and charging.
@@
Figure 6: The framework of the proposed SEAL defense scheme in UAV offloading scenario [160].
@@ 5.2 Physical-Layer security
Physical layer security of LAWNs refers to protecting the confidentiality, integrity, and availability of information transmitted and processed within LAWNs by exploiting the inherent characteristics of the physical layer itself [153, 167]. This protection is achieved without relying solely on traditional cryptographic methods. By addressing the physical layer security concerns of the above issues, LAWNs can enhance resilience against sophisticated attacks, enable robust operation in contested environments, and provide lightweight security solutions suitable for resource-constrained UAV platforms, ultimately contributing to safer and more reliable UAV deployments. In general, physical layer security is commonly classified into two fundamental paradigms, namely the wiretap channel and covert communication. Currently, researchers have adopted beamforming techniques, artificial noise injection, cooperative jamming, power allocation strategies, and intelligent reflecting surfaces (IRS) to realize the physical layer security of LAWNs on sensing, communication, computing, and control aspects. For example, for physical layer security in integrated sensing and communication, [176] explored a system where a UAV transmits ISAC waveforms to securely communicate with IoT devices while simultaneously sensing the environment. This approach optimizes UAV trajectory and beamforming to maximize the average communication secrecy rate while ensuring accurate sensing capabilities. In the context of covert ISAC in LAWN systems, [174] studied covert ISAC for multi-antenna UAV systems using a unified waveform for communication and sensing, revealing a trade-off between covert communication and sensing accuracy by exploiting prior information of the adversary. Regarding physical layer security in UAV communication, [90] addressed wire-tap channel communications in 5G networks with eavesdroppers by proposing a virtual antenna array with collaborative beamforming. To enhance security, UAV positions and excitation current weights are optimized, aiming to improve total secrecy rates and reduce sidelobe levels through an improved multi-objective dragonfly algorithm. Furthermore, to satisfy the heterogeneous security requirements in low-altitude UAV systems, where some users demand secrecy while others require covertness, the authors in [60] investigated multi-user cooperative secret and covert communications for UAV networks. For the physical layer security aspect within computing, [210] investigated a secure UAV MEC system where ground users offload tasks to a legitimate UAV amidst eavesdropping UAVs. It enhances security by employing jamming signals from the legitimate UAV and users to maximize minimum secrecy capacity, jointly optimizing aspects like offloading ratio and computing capacity to balance security and latency. [122] focused on secure and energy-efficient UAV-MEC with RIS assistance, aiming to maximize minimum energy efficiency against an eavesdropper. The research ensured security through artificial noise transmission and optimizes computation tasks/CPU frequency allocation alongside other parameters, demonstrating strong performance even with an eavesdropper present. Focusing on physical layer security in control, [178] proposed a secret and covert transmission strategy for ground-to-UAV links using non-orthogonal multiple access (NOMA). This method allows simultaneous transmissions with different security needs, enabling an adversary to attempt both data eavesdropping and transmission detection. Security is rigorously ensured with optimization targeting maximized rates through beamforming and trajectory control. [34] improved covert communication range by using a UAV relay and finite blocklength transmissions, aiming to maximize secure bits against a warden. It achieves this by optimizing the warden’s threat potential and jointly controlling transmission parameters under error constraints, leveraging signal uncertainty at the warden for physical layer covertness.
@@ 5.3 Network-Layer security
For example, addressing network layer security challenges for UAV-based sensing, [141] designed an intrusion detection and response system to identify and mitigate attacks that compromise the integrity of sensed data, including false information dissemination and GPS spoofing. To ensure network layer security in UAV communication, [52] introduced self-adaptive intrusion detection (SID)-UAV, which employs a self-matching system to detect safe communication routes, coupled with a distributed architecture for investigation and response to malicious UAVs. The work of [111] enhanced network layer security for autonomous aerial vehicle (AAV)-aided short-packet transmissions against detection and eavesdropping by optimizing for maximum covert secrecy energy efficiency through a dual-loop iterative algorithm for user association, power control, and trajectory planning. Addressing network layer security in UAV edge computing, [140] presented a cyber-defense mechanism based on a non-cooperative game to protect against network and offloading threats. The solution aims to maximize protection while minimizing the computational and energy costs associated with the defense. In response to the data leakage risk in UAV-assisted traffic surveillance at the network layer, [56] utilized a probabilistic data structure approach for cyber-threat detection and efficient data management, minimizing computational overhead in the process. Regarding network layer security for control, [11] proposed using sliding mode control to enhance UAV security. The system dynamically reconfigures to adapt to network-level threats like GPS spoofing and jamming, ensuring robust control even under attack through anomaly detection and fault tolerance. [172] tackled network layer access control challenges in distributed UAV swarm control by using a blockchain-based scheme. It ensures secure authentication and data integrity against identity spoofing and data injection, employing dynamic key management for task and identity groups.
@@ 6.1 Large AI Model-based LAWNs
Large AI models, characterized by their extensive parameter dimensions and training on massive datasets, have made notable advancements across various domains, including natural language processing, computer vision, and autonomous systems. Prominent models, e.g., generative pre-trained transformers (GPT), bidirectional encoder representations from transformers (BERT), and similar architectures, possess the ability to learn intricate patterns and generalize across a wide range of tasks. This makes them particularly suitable for dynamic and complex environments. Hence, the integration of large AI models into LAWNs holds substantial potential for enhancing the cognitive capabilities of UAVs, facilitating real-time decision-making in unpredictable environments. A common approach to this integration involves utilizing large language models (LLMs) in physical layer designs, where pre-trained models are deployed at both the UAVs and controller terminal to enable effective online inference. For example, the work of [171] illustrates how large AI models can be employed in UAV path planning, where they enable the UAVs to dynamically adjust their flight paths based on real-time environmental data, such as wind and obstacle conditions. Similarly, the work of [70] explores how large language models (LLMs) can enhance UAV communication by enabling more intuitive interactions between UAVs and ground control stations. This capability allows for more adaptive and flexible communication, improving operational coordination, especially in complex scenarios like disaster relief or surveillance operations.
@@
Model distribution involves partitioning a large AI model across multiple devices or computational nodes, such as UAVs or edge servers. Instead of running the entire model on a single device, each UAV or edge server processes a portion of the model. This distributed approach allows for parallel processing, which significantly reduces the computational burden on each individual device. The work of [127] demonstrated how model distribution can be effectively implemented in UAV networks, where large AI models are divided across several UAVs, enabling real-time inference without overwhelming the resources of any single UAV. This technique enhances the scalability of large models, making it possible to deploy them in a distributed manner without compromising system performance. Data distribution, on the other hand, focuses on distributing the data processing tasks across multiple devices to minimize the need for large-scale data transmission. In the context of LAWNs, this approach ensures that data collected by UAVs is processed locally, reducing the communication overhead and enabling real-time decision-making. The work of [70] highlighted the use of federated learning in UAV networks, a data distribution method that allows multiple UAVs to collaboratively train a shared model by exchanging model updates rather than raw data. This approach not only preserves data privacy but also reduces the bandwidth required for communication between UAVs and ground control systems, making it particularly beneficial in applications where security and privacy are paramount.
@@
However, it is important to note that both model distribution and data distribution entail substantial resource-intensive fine-tuning processes [100]. Therefore, high-dimensional computing (HDC) has attracted increasing attention as a computationally efficient alternative [29]. HDC is a brain-inspired computing paradigm in which information is encoded into extremely high-dimensional vectors, often containing tens of thousands of components. Such representations naturally support large-scale parallelism and enable rapid processing, while exhibiting strong robustness against noise, distortions, and missing data. A fundamental property of these high-dimensional vectors is that independently generated vectors are almost orthogonal to one another. This behavior results from the randomization procedures during vector construction, which ensure that distinct vectors share very limited overlap [18]. Consequently, meaningful representations tend to have large Hamming distances and low cosine similarity, providing substantial capacity for encoding diverse information with minimal interference. These characteristics enable HDC to rely on simple arithmetic operations, including addition for combining multiple vectors and multiplication for binding variables with their associated values [80]. Similarity measures, such as Hamming distance and cosine similarity, are used to determine how closely two vectors align, supporting accurate retrieval and comparison. Since these operations do not require heavy matrix computations, they are highly efficient and particularly suitable for UAVs operating in LAWNs, where memory, computation capability, and energy supply are inherently limited.
@@
HDC has also shown strong potential for improving wireless communication reliability in scenarios relevant to LAWNs. Its high-dimensional representations can tolerate interference and signal distortions, allowing communication to remain stable even when multiple devices transmit simultaneously. A representative example is the direct transmission of high-dimensional vectors from a group of sensing devices. Instead of coordinating access or assigning dedicated channels, each device simply sends the vector that corresponds to its measurement. The receiver then observes the superimposed vector and determines the underlying sensing events by comparing the received vector with a set of reference patterns. By doing so, HDC enables efficient UAV reporting without traditional access control. For example, the authors in [69] considered a scenario where the receiver detects temperature conditions and determines how many devices experienced a particular temperature by analyzing the combined vector. Furthermore, the study in [83] generated composite vectors through binding and superposition to encode local sensing information at each device. The receiver successfully interpreted the embedded information using predefined base vectors, demonstrating that HDC can maintain reliable information extraction even in dense and interference-prone environments. These results indicate that HDC offers a practical and lightweight computational method for LAWNs, enabling robust communication and efficient information exchange while accommodating the limitations of UAV platforms.